A researcher wishes to estimate the proportion of adults who have​ high-speed Internet access. What size sample should be obtained if she wishes the estimate to be within 0.04 with 90​% confidence if
​(a) she uses a previous estimate of 0.58​?
​(b) she does not use any prior​ estimates?

Answer by rothauserc(4718)

Margin of Error(ME) = 0.04
:
critical statistic(z) for a 90% confidence level is 1.645
:
when standard deviation is not known, we use
:
sample size n = p*(1-p)*(z/ME)^2, where p is the proportion
:
a) n = 0.58 * (1-0.58)*(1.645/0.04)^2 = 411.9923
:
sample size is 412
:
b) p is not known, so use p = 0.50
:
n = 0.50 * (1-0.50)*(1.645/0.04)^2 = 422.8164
:
sample size is 423
: